• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 23:03:55 on October17,2010

    Entry number 335406


    Compton background rates have been very bad(16000Hz/uA) since beam is back. I stopped the DAQ. I also turned off the Cavity and Photon detector as well.

    FIGURE 1